this is on a test of mine that i am having difficulty with, this is the ENTIRE QUESTION on the test in front of me, i am not cutting the question out on purpose.
Of the fifty ballots cast in the last election, at least one of them had an irregularity.

Found 2 solutions by ikleyn, rothauserc:
Answer by ikleyn(52781)

.

This set of words does not contain a question.

Any set of words with no question IS NOT a math problem.

COROLLARY. The set of words in your post IS NOT a Math problem.



Answer by rothauserc(4718)

Probability(P) at least one of them had an irregularity = 1 - P(that none of the 50 ballots had an irregularity)
:
This is an application of the Binomial Probability Formula,
:
P(k successes out of n trials) = nCk * p^k * (1-p)^(n-k), where nCk = n!/(k! * (n-k)!), k is number of successes, n is the number of trials
:
If the problem does not give you the probability of a ballot having an irregularity, then you have to use 50%, that is, p = 0.50
:
n = 50 and k = 0 means that 50C0 = 1 and p^0 = 1 but still need p to calculate 1 - p
: